OpenStudy (lgbasallote):

The mineral pyromorphite has the formula \(\text{Pb}_5(\text{PO}_4)_3\text{Cl}\). What mass percent of chlorine does it contain? i use \[\text{mass %} = \frac{\text{mass of Cl}}{\text{mass of} \;\text{Pb}_5(\text{PO}_4)_3\text{Cl}}\times 100\text %\] right?

OpenStudy (lgbasallote):

i got \[\frac{35.45}{1,356.36} \times 100\text % \implies 0.0261 \times 100 \text %\implies 2.61 \text%\] is that right?

OpenStudy (australopithecus):

Pb5(PO4)3Cl Pick a random mass of Pb5(PO4)3Cl to work with I'm going to pick 100g \[(\frac{( 100g_{Pb_5(PO_4)_3Cl} *\frac{35.45g/mol_{Cl}}{1356g/mol_{Pb_5(PO_4)_3Cl}} )}{100g})* 100 =2.6 \] Your method looks good I just did it the longer way molecular mass ratio is constant so I didnt really need to apply gravimeter factor :)

OpenStudy (australopithecus):

sorry that should be 2.61% :)

OpenStudy (lgbasallote):

the answer says it shoulld be 10.50% @Australopithecus that's why i was asking

OpenStudy (australopithecus):

I think your source is wrong

OpenStudy (lgbasallote):

i doubt that

OpenStudy (australopithecus):

hmm well our answer makes sense logically as well chlorine takes up only a small fraction of the mass that all the other elements take up, I mean it has 5 atoms of lead in it

OpenStudy (australopithecus):

and only 1 of chlorine, what makes you doubt our answer?

OpenStudy (lgbasallote):

what makes me doubt is im not good in chemistry

OpenStudy (australopithecus):

Well I'm a laboratory technologist and currently obtaining a degree in biochemistry, trust me it is the correct answer

OpenStudy (australopithecus):

The general equation should be, \[(\frac{Molecular Mass_{Cl}}{Molecular Mass_{Pb_5(PO_4)_3Cl}})*100 = Mass_{Cl}%\]

OpenStudy (lgbasallote):

well i'll still want to see how 10.5 was in the choices

OpenStudy (australopithecus):

That should be mass percent Cl

OpenStudy (lgbasallote):

people dont post choices unless they make sense

OpenStudy (australopithecus):

Must have been a miss calculation of the molecular mass of pyromorphite

OpenStudy (australopithecus):

I used wolframalpha to check its molecular mass btw

OpenStudy (lgbasallote):

then how to get that miscalculation?

OpenStudy (australopithecus):

https://www.wolframalpha.com/input/?i=+Pb5%28PO4%293Cl

OpenStudy (australopithecus):

who knows, but we both came to the same conclusion and furthermore our solution makes sense 10% doesn't make any sense look at the compound the majority of mass will be taken up by lead

OpenStudy (lgbasallote):

i think oxygen takes the bigger percent...

OpenStudy (australopithecus):

no, do you understand what molecular mass is?

OpenStudy (australopithecus):

it is the ratio of mass per mole Oxygens molecular mass is approximately (15.9994g/mol )*4 = (63.9976g/mol)*3 = 191.9928g/mol Lets look at leads molecular weight in this compound 207.2g/mol = Molecular mass of Lead 207.2g/mol*5 = 1036g/mol of lead In the compound

OpenStudy (lgbasallote):

it seems from this website it really is 2.61 http://webmineral.com/data/Pyromorphite.shtml

OpenStudy (lgbasallote):

i think i do understand the concept of molecular mass btw....i did answer this...lol

OpenStudy (australopithecus):

chlorine has 35.45grams per 1mol of compound lead has 1035g per 1 mol of compound

OpenStudy (australopithecus):

Well you should have known you were right logically, and furthermore you stated that oxygen takes up more mass than lead in the compound :S

OpenStudy (lgbasallote):

lol no need to explain. i already showed that i got 2.61...what i was asking how 10.5 was got

OpenStudy (australopithecus):

That is a mystery that is not worth solving :)

OpenStudy (lgbasallote):

i was referring to the mole ratio that oxygen takes more

OpenStudy (australopithecus):

lead has 1035g per 1 mol of compound Oxygen has 191.9928g per 1 mol of compound

OpenStudy (australopithecus):

it is an incorrect assumption on your part I just wanted to point out

OpenStudy (australopithecus):

I'm not trying to be a jerk btw :)

OpenStudy (australopithecus):

you always have to remember molecular mass is expressed per 1 mol Cl has a molecular mass of \[\frac{35.45g}{1mol}\]

OpenStudy (lgbasallote):

i was lol

OpenStudy (australopithecus):

I'm sorry you stated you are not good at chemistry I just wanted to make sure you knew please don't take offense